A perfect competitive firm has a total revenue and total cost curves given as.TC=100Q and TC =5000+2Q+0.2Q^2 find the profit maximising out of.the firm. What profit does the.firm make?
1
Expert's answer
2018-05-01T09:47:08-0400
A perfect competitive firm has a total revenue and total cost curves given as: TR = 100Q and TC = 5000 + 2Q + 0.2Q^2.
The profit-maximising output of the firm is at MR = MC = P, then:
TR' = TC',
100 = 2 + 0.4Q,
Q = 245 units,
P = $100.
The profit is:
TP = TR - TC = 100*245 - (5000 + 2*245 + 0.2*245^2) = $7005.